1. On June 30, 2021, Hercule, Inc. leased warehouse equipment from Marble, Inc. The lease agreement...

1.

On June 30, 2021, Hercule, Inc. leased warehouse equipment from Marble, Inc. The lease agreement calls for Hercule to make semiannual lease payments of $843,755 over a three-year lease term, payable each June 30 and December 31, with the first payment at June 30, 2021. Hercule’s incremental borrowing rate is 8%, the same rate Marble used to calculate lease payment amounts. Marble manufactured the equipment at a cost of $4.0 million. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
1. Determine the price at which Marble is “selling” the equipment (present value of the lease payments) at June 30, 2021.
2. What amounts related to the lease would Marble report in its balance sheet at December 31, 2021? (Ignore taxes.)
3. What would be the net effect of the lease that Marble would report in its income statement for the year ended December 31, 2021? (Ignore taxes.)

2.

I. Lasch Co. recorded a right-of-use asset of $220,000 in a 10-year operating lease. Payments of $37,356 are made annually at the end of each year. The interest rate charged by the lessor was 11%. The balance in the right-of-use asset after the first year will be:

195,800

206,844

142,514

220,000
Answer #1

You have asked multiple unrelated questions in the same post. As if they were not enough, one of your questions has multiple sub parts. I have therefore addressed all the sub parts of the first question. Please post the balance questions separately.

1. The price at which Marble is “selling” the equipment = present value of the lease payments at June 30, 2021 = Lease payment x PVAD (i = 8%/2 =4%, n = 3 x 2 = 6) = 843,755 x 5.4518 = $ 4,599,984

Part 2 & 3. First lease payment is on the first day of lease i.e. as soon as lease is entered into. Hence, the entire payment will be used towards reduction of lease receivable as there is no interest accrued. Hence, lease payable balance after the first payment on June 30 = 4,000,000 - 843,755 = 3,156,245

Interest for 6 months =  3,156,245 x 8% x 6/12 = 126,250

Payment towards principal = 843,755 - 126,250 =  717,505

Hence, receivable balance at the end of December 31, 2021 = 3,156,245 -  717,505 =  2,438,740

Hence, your final answers are:

Part 2 On the balance sheet Pretax amount of net receivable

  2,438,740

Part 3. On the income statement Pretax amount of interest revenue 126,250